OpenStudy (anonymous):

Given that f(x)=4x^2, and g(x)= 3-(x/2), what is the value of f(g(4))?

OpenStudy (anonymous):

you need to find \(g(4)\) first did you do that?

OpenStudy (anonymous):

i haven't been taught that yet.

OpenStudy (anonymous):

\[g(x)=3-\frac{x}{2}\] to find \(g(4)\) replace \(x\) by \(4\) and compute \[g(4)=3-\frac{4}{2}\] what do you get?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

ok good now you need \(f(g(4))=f(1)\)

OpenStudy (anonymous):

thanx. ik it sounds like a dumb question. :(

OpenStudy (anonymous):

\(f(x)=4x^2\) \(f(1)=?\)

OpenStudy (anonymous):

16

OpenStudy (anonymous):

oh careful here! exponents first, multiplication second

OpenStudy (anonymous):

\[f(1)=4\times 1^2=4\times 1=4\] not \(16\)

OpenStudy (anonymous):

thanx. ik it sounds like a dumb question. :(